OpenStudy (anonymous):

Name the Property of Congruence that justifies the statement: If .MN is congruent to LK then LK is congruent to MN A. Symmetric Property C. Reflexive Property D. none of these

OpenStudy (anonymous):

i think the answer is A

OpenStudy (mertsj):

You are correct. The symmetric property of equality states: If a = b then b = a

OpenStudy (mertsj):

And the symmetric property of congruence is the same thing.

OpenStudy (anonymous):

Complete the two-column proof.

OpenStudy (anonymous):

Given: x/5+9=11 Prove: x=10

OpenStudy (anonymous):

x/5+9=10 a________________? x/5=2 b_____________? x=10 c_____________?

OpenStudy (anonymous):

A. a. Given b. Subtraction Property of Equality c. Division Property of Equality B. a. Given b. Subtraction Property of Equality c. Multiplication Property of Equality D. a. Given b. Addition Property of Equality c. Division Property of Equality

OpenStudy (anonymous):

i think b is the correct answer

OpenStudy (mertsj):

B is correct

OpenStudy (mertsj):

How did you miss these if you know all the right answers?

OpenStudy (anonymous):

i just want to confirm the please can you help because i still believe some of are still wrong

OpenStudy (anonymous):

http://prntscr.com/ymi1q

OpenStudy (anonymous):

can you see the pic and tell me which are the alternate angles

OpenStudy (mertsj):

Do you mean alternate interior angles or alternate exterior angles?

OpenStudy (anonymous):

A: <FEG and <CBE are alternate angles. C: <CBE and <DEB are same-side interior angles. D: <FEG and <DEB are same-side interior angles.

OpenStudy (anonymous):

i think the answer is A

OpenStudy (mertsj):

C

OpenStudy (anonymous):

http://prntscr.com/ymiv7

OpenStudy (anonymous):

option D is none of these

OpenStudy (anonymous):

the question is in the link

OpenStudy (mertsj):

A

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

http://prntscr.com/ymjpi

OpenStudy (anonymous):

http://prntscr.com/ymjv9

OpenStudy (anonymous):

question in the link

OpenStudy (mertsj):

A

OpenStudy (anonymous):

i think the answer is A

OpenStudy (mertsj):

I agree

OpenStudy (anonymous):

http://prntscr.com/ymk5p

OpenStudy (mertsj):

Are there only two choices?

OpenStudy (anonymous):

http://prntscr.com/ymkhv

OpenStudy (anonymous):

i think the answer is D

OpenStudy (anonymous):

question is in the link

OpenStudy (mertsj):

I can't tell the difference between B and C

OpenStudy (mertsj):

And I think it should be like this:

OpenStudy (anonymous):

thats the problem iam facing

OpenStudy (anonymous):

so i opted out D

OpenStudy (anonymous):

what are your views

OpenStudy (mertsj):